# Distributionally Robust Markov Games with Average Reward

### v2

@@ D.2 Hemi-continuity
Let $\{\pi^{n}_{-i}\}_{n\geq 1}\subset\prod_{s}\Delta(\mathcal{A}_{-i})$ satisfy $\pi^{n}_{-i}\to\pi_{-i}$ (product topology). For each $n$, let $(u_{n},\alpha_{n})$ solve the attained sup–inf constant-gain equation
@@
converges to $(\mathcal{P}^{\pi_{-i}})^{s}_{a_{i}}$ in the Hausdorff metric because the right-hand side is a Minkowski sum with convex weights that depend continuously on $\pi_{-i}(\cdot\mid s)$; hence $(\mathcal{P}^{\pi^{n}_{-i}})_{s}\to(\mathcal{P}^{\pi_{-i}})_{s}$ in the Hausdorff metric (Lemma˜D.3).
@@ D.3 Existence of DR-NE
We can then prove the existence of DR-NE following the results shown above. Define a map $G(\pi)\triangleq\bigotimes_{i\in\mathcal{N}}\mathcal{G}_{i}(\pi_{-i})$ on the compact convex policy space $X=\prod_{i,s}\Delta(\mathcal{A}_{i})\to 2^{X}$. Since each $\mathcal{G}_{i}(\cdot)$ is nonempty, compact, convex (Theorem˜D.1), and is upper hemi-continuous (Theorem˜D.2), thus $G$ satisfies Kakutani’s condition as in Lemma˜4.2. Hence there exists a fixed point $\pi^{\star}\in G(\pi^{\star})\subseteq\bigotimes_{i}\text{BR}_{i}(\pi^{\star}_{-i})$, which is a stationary DR-NE.
@@ Appendix E Proof of Section 6
Such an assumption is necessary for convergence (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence).
@@
We first show that, if the algorithm converges to $\pi$, then $\pi$ is an average reward Nash Equilibrium. To show this, we note that the convergence of Alg 1 implies that there exists some constant $c_{1},...,c_{N}$, such that
$\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i})]-h_{i}(s)=c_{i},\forall i,s.$ (92)
This implies that
$h_{i}(s)+c_{i}=\sum_{a}\pi(a|s)(r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})),$ (93)
which is the robust Bellman equation w.r.t. $\pi$. Hence by Theorem 4.1, $(c_{i},h_{i})=(g^{\pi}_{\mathcal{P},i},h^{\pi}_{\mathcal{P},i})$ for some vector $h^{\pi}_{\mathcal{P},i}$, i.e.,
$h^{\pi}_{\mathcal{P},i}+g^{\pi}_{\mathcal{P},i}=r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i}).$ (94)

### v3

@@ D.2 Hemi-continuity
Let $\{\pi^{n}_{-i}\}_{n\geq 1}\subset\times_{s}\Delta(\mathcal{A}_{-i})$ satisfy $\pi^{n}_{-i}\to\pi_{-i}$ (product topology). For each $n$, let $(u_{n},\alpha_{n})$ solve the attained sup–inf constant-gain equation
@@
converges to $(\mathcal{P}^{\pi_{-i}})^{s}_{a_{i}}$ in the Hausdorff metric because the right-hand side is a Minkowski sum with convex weights that depend continuously on $\pi_{-i}(\cdot\mid s)$; hence $(\mathcal{P}^{\pi^{n}_{-i}})_{s}\to(\mathcal{P}^{\pi_{-i}})_{s}$ in the Hausdorff metric (Lemma˜D.4).
@@ D.3 Existence of DR-NE
We can then prove the existence of DR-NE following the results shown above. Define a map $\mathcal{G}(\pi)\triangleq\times_{i\in\mathcal{N}}\mathcal{G}_{i}(\pi_{-i})$ on the compact convex policy space $X=\times_{i,s}\Delta(\mathcal{A}_{i})\to 2^{X}$. Since each $\mathcal{G}_{i}(\cdot)$ is nonempty, compact, convex (Theorem˜D.1), and is upper hemi-continuous (Theorem˜D.2), thus $G$ satisfies Kakutani’s condition as in Lemma˜4.2. Hence there exists a fixed point $\pi^{\star}\in\mathcal{G}(\pi^{\star})\subseteq\times_{i}\text{BR}_{i}(\pi^{\star}_{-i})$, which is a stationary DR-NE.
@@ E.1 Nash Iteration
Moreover, we assume that $P^{a}_{s}>0$ for all $s,a$ and $P\in\mathcal{P}^{a}_{s}$.
The first part of our assumption is necessary for convergence under average reward Markov games (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). The second assumption is a technical assumption to ensure convergence, and can be removed by modifying the update rule by a multi-step operator, see (wang2023model; xu2025finite).
@@
We split the proof into two parts.
(i) Any limit point is an average-reward NE.
Suppose Algorithm 1 converges to a stationary joint policy $\pi$, and let $h_{i}$ be the limit of the bias estimates for player $i$. By the update rule in Algorithm 1, at iteration $k$,
$h_{i}^{k+1}(s)=\mathbb{E}_{\pi^{k}(s)}\bigl[Q_{i}(s,a,h_{i}^{k})\bigr],$
where $\pi^{k}(s)$ is the selected NE of the stage game with payoff matrix $Q_{i}(s,a,h_{i}^{k})$. At convergence we have $h_{i}^{k+1}\to h_{i}^{*}$ and $\pi^{k}\to\pi$, and the stopping condition $\max_{i}\operatorname{sp}(h_{i}^{k+1}-h_{i}^{k})=0$ implies that the difference $h_{i}^{k+1}-h_{i}^{k}$ converges to a constant vector $c_{i}\mathbf{1}$ for each $i$ (here $\operatorname{sp}(\cdot)$ denotes the span seminorm). Thus there exist constants $c_{1},\dots,c_{N}$ such that for all $s$ and $i$,
$\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i}^{*})]-h_{i}^{*}(s)=c_{i}.$ (87)
Unfolding the definition of $Q_{i}$,
$Q_{i}(s,a,h_{i}^{*})=r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i}^{*}),$
and taking the expectation in (95) gives
$h_{i}^{*}(s)+c_{i}=\sum_{a}\pi(a\mid s)\Bigl(r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i}^{*})\Bigr),\quad s\in\mathcal{S}.$ (88)
Equation (88) is the robust average-reward Bellman equation associated with the fixed policy $\pi$ and bias $h_{i}^{*}$. By the solvability result for the robust Bellman equation under Assumption 1 (Theorem 4.1 in the main text, based on the single-agent theory in wang2025bellman), it follows that $(c_{i},h_{i}^{*})=(g^{\pi}_{P,i},h^{\pi}_{P,i})$ for some robust bias $h^{\pi}_{P,i}$, i.e.,
$h^{\pi}_{P,i}+g^{\pi}_{P,i}\mathbf{1}=r^{\pi}_{i}+\sigma_{\pi}(h^{\pi}_{P,i}).$ (89)